OpenStudy (anonymous):

A building in the city has a rectangular base. The length of the bae measures 70 feet less than 3 times the width. The perimeter of the bae is 860 ft. What are the dimensions of the base?

OpenStudy (anonymous):

@satellite73 @jim_thompson5910

OpenStudy (anonymous):

please I really need help

OpenStudy (midhun.madhu1987):

Let the width be x so length = 70 - 3x Perimeter = 860 2(l + b) = 860

OpenStudy (midhun.madhu1987):

just plug in the values of l and b and solve for x